OpenStudy (anonymous):

what is the remainder when 4^31 is divided by 10

OpenStudy (kewlgeek555):

Hold on, sorry. Let me try and see if I can solve this..

OpenStudy (kewlgeek555):

Answer: 4 This can be done with or without a calculator. Calculator method: Punch it into a calculator and see what the remainder is. Without calculator: Investigate power of 4 and the remainder when it is divided by 10. 4^1 = 4 (the remainder is 4) 4^2 = 16 (the remainder is 6) 4^3 = 64 (the remainder is 4) 4^4 = 256 (the remainder is 6) This pattern continues. If the power is odd, then you will get a remainder of 4.